OpenStudy (anonymous):

Can anyone help me with this? Determine the angle between the following vectors in R4: (i) v1 = (1, 0, 2, 2) and v2 = (4, 2, 1, 2);

OpenStudy (anonymous):

are you asking me

OpenStudy (anonymous):

no but i can try use what i have in vector analysis

OpenStudy (calculusfunctions):

If two vectors u and v have an angle θ between them, then\[\cos \theta =\frac{ u∙v }{ \left| u \right|\left| v \right| }\]

OpenStudy (calculusfunctions):

@Sportaholic013 understand?

OpenStudy (anonymous):

Thank you :) Ya I get it now :D

OpenStudy (calculusfunctions):

Great!